You want to put $8,000 in a simple interest account. It has a 2% annual interest rate. How long must you invest that money to earn $400 in interest? Use I=prt

Answer by mananth(16949)

I = P * r * t
P= 8000
r=2% = 0.02
t=?
I = 400
400=8000*0.02*t
400/(8000*0.02) =t
400/160=t
2.5 = t
2 1/2 years